1. The sum of two numbers is 24. The second number is 3 times the first number.
a. Write a system of equations that represents the information in
b. Solve the system of equations using the Elimination Method
c. Solve the system of equations using the Substitution Method

Answer by mananth(16949)

The sum of two numbers is 24. The second number is 3 times the first number.
a. Write a system of equations that represents the information in
b. Solve the system of equations using the Elimination Method
c. Solve the system of equations using the Substitution Method
Let one number be x
The other number will be 24-x
24-x = 3x
24 = 4x
24/4 =x
x=6 & 18 is the other number
------------------------------
Let one number be x
the other number be y
y=3x
x+y=24 Substitute the value of y in the equation
x+3x=24
4x=24
x=6 y=3x so y=18
--------------------------
x+y =24 --------------1
y=3x
y-3x=0 ---------------2
subtract equartion 2 from 1
x+y- (y-3x)=24
x+y-y+3x=24
x+3x=24
4x=24
x=6
y=3x
y=18
The numbers are 6 & 18